(a) Derive the demand functions for the utility function U=(a)sqrt(x)+(b)sqrt(z) +xz

(b) Let a = 2, b = 3, px = 1, pz = 2, and Y = 50. Find the optimal values for x and z.
Answer #1

J. Uz a? x + b3z + x2 Mox - a² + 2 MUY = b²+x MRS = a²+2 b²+x Platz) = Px (6²7) Pza²+ P2.2 = Px (6²+X) Pz. 2 = Px (6²+x) - Pez fes (0+ yerda Pao) - (6) het a= 2, b=3 : Px=1 ; P2=2 : Y = 50 2- 4(3= (23** [.50 - 1639) 3-44 (5048-1) 18-16 + 4g 2- X= - [